# DAOcc: 3D Object Detection Assisted Multi-Sensor Fusion for 3D Occupancy Prediction

### v1

@@ Abstract
Multi-sensor fusion significantly enhances the accuracy and robustness of 3D semantic occupancy prediction, which is crucial for autonomous driving and robotics. However, existing approaches depend on large image resolutions and complex networks to achieve top performance, hindering their application in practical scenarios. Additionally, most multi-sensor fusion approaches focus on improving fusion features while overlooking the exploration of supervision strategies for these features. To this end, we propose DAOcc, a novel multi-sensor fusion occupancy network that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image feature extraction network and practical input image resolution. Furthermore, we introduce a BEV View Range Extension strategy to mitigate the adverse effects of reduced image resolution. As a result, our approach achieves new state-of-the-art results on the Occ3D-nuScenes and SurroundOcc datasets, using ResNet50 and a 256x704 input image resolution. Code will be made available at https://github.com/AlphaPlusTT/DAOcc.
@@ 1 Introduction
Figure 1: Unlike existing multi-modal methods [46, 33, 58, 40, 34], which rely on complex combinations of loss functions, including several or all of the focal loss $\mathcal{L}_{focal}$ [38], the scene-class affinity loss $\mathcal{L}_{scal}^{geo}$ and $\mathcal{L}_{scal}^{sem}$ [4], the lovasz-softmax loss $\mathcal{L}_{ls}$ [2], and the depth loss $\mathcal{L}_{d}$ [21], to achieve optimal performance, our approach only requires a single cross-entropy loss $\mathcal{L}_{ce}$ for the occupancy prediction task.
@@
In existing works on multi-modal [57, 33, 58, 49, 40, 46, 34] or image-based [4, 48, 41, 12, 47, 23, 54, 39, 50] occ tasks, achieving superior performance often involves using extremely high image resolutions and complex image feature extraction networks (see Table 1 and Table 3), such as using 900×1600 resolution input image and ResNet101 [11] equipped with DCN [7, 62]. However, this approach significantly limits the deployment of top-performing occ models on edge devices due to the high computational demands. Compared to image, point cloud is much sparser. For example, in the training set of the nuScenes dataset, the maximum number of points in a single frame point cloud is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image. Therefore, how to effectively leverage point cloud data within a multi-modal occ framework remains to be further explored.
@@
Moreover, we observe that most works [57, 33, 58, 49, 40, 46] on multi-modal occ primarily focuses on obtaining more effective fusion features, with insufficient exploration into the forms of supervision for these fused features. Although CO-Occ [34] introduces regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometric and structural information inherent in point cloud features. In contrast, point cloud-based 3D object detectors [56, 6, 31] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based detectors [26, 16, 27, 28]. This observation suggests a new research direction: how to better exploit the unique strengths of point cloud data in multi-modal occ tasks.
@@
Figure 2: Overview of our proposed DAOcc. We first introduce the BVRE strategy to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective. Then feature extraction is performed through the multi-modal occupancy prediction network illustrated in the above figure. Furthermore, to fully leverage the inherent geometric structures within point cloud features, we incorporate 3D object detection as auxiliary supervision.
@@
Based on these observations, we propose DAOcc, a novel multi-modal occ framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image feature extraction network and practical input image resolution.
@@
In constructing the baseline network structure of DAOcc, we adopt the most direct and straightforward approach: first, we extract features from images and point clouds, respectively. Given that depth estimation from monocular images is an ill-posed problem [13, 20] and that deformable attention modules are overly complex [10], we employ a simple method, similar to Harley et al. [9], to transform image features from 2D space to 3D volumetric space. Specifically, we project a set of predefined 3D points onto the 2D image feature plane and use bilinear interpolation to sample the corresponding 2D image features for these 3D points. Next, we adopt the same simplest fusion strategy as BEVFusion. [30], concatenating image and point cloud features, then performing a 2D convolution to fuse them, leading to a unified BEV features. Finally, we apply a fully convolutional BEV encoder with a residual structure to further fuse the unified BEV features, and subsequently employ the Channel-to-Height operation [54] to reshape the height of the unified BEV features.
@@
To fully exploit the geometric and structural information inherent in point cloud features, we augment the baseline model’s unified BEV features with 3D object detection supervision, thereby enhancing the discriminability of the unified BEV features. This renders the unified BEV features more sensitive to object boundaries and capable of perceiving relationships between internal object structures. Additionally, given the sparsity of point clouds, we extend the processing range of the point clouds and employ sparse convolutions [52] to mitigate the computational overhead introduced by this extension. We refer to this approach as BVRE (BEV View Range Extension). BVRE provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution. It is worth noting that the 3D object detection supervision is only used as an auxiliary branch during training and can be removed during inference for the occ task.
@@
As a result, the proposed DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes and SurroundOcc benchmarks, while using ResNet50 and a 256×704 input image resolution. Specifically, on the Occ3D-nuScenes validation set, DAOcc achieves 53.82 mIoU when using the camera mask during training and 48.2 RayIoU without using the camera mask. Additionally, on the SurroundOcc validation set, DAOcc achieves 45.0 IoU and 30.5 mIoU.
@@
  • We design a simple yet efficient multi-modal 3D semantic occupancy prediction baseline, eliminating the need for complex deformable attention modules as well as image depth estimation during feature fusion.
  • We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image feature extraction network and input image resolution.
@@
  • We establish the new state-of-the-art performance on Occ3D-nuScenes dataset and SurroundOcc dataset, while using ResNet50 and a 256×704 input image resolution.
@@ 2 Related Work
Camera-based 3D occupancy prediction has gained significant attention due to its cost-effectiveness. MonoScene [4] is the first work to infer the dense occupancy and semantics of outdoor and indoor scenes from a single RGB image. Since full perception of the surrounding environment is crucial for making accurate decisions in autonomous driving and robotics, most of the recent work is based on surround view image input. TPVFormer [15] proposes an efficient tri-perspective view (TPV) representation that combines BEV with two additional perpendicular planes to provide a 3D perception result with multi-view image inputs. Given the insufficient fine-grained semantic information of the TPV representation, OccFormer [59] leverages dense 3D features and proposes a dual-path transformer-based occupancy network. However, both TPVFormer and OccFormer use sparse LiDAR points as supervision, resulting in sparse occupancy predictions. To obtain dense occupancy prediction, OpenOccupancy [46], SurroundOcc [48] and Occ3D [41] developed methods for generating dense occupancy labels and established benchmarks on their respective proposed datasets. It is worth noting that SurroundOcc does not introduce manual annotation. The emergence of these benchmarks further promoted the development of 3D occupancy prediction. PanoOcc [47] adopts voxel query and a coarse-to-fine approach to learn a unified occupancy representation. FlashOcc [54] uses BEV features with 2D convolution to achieve efficient feature extraction and proposes the channel-to-height transformation to lift the output of BEV to 3D space. SparseOcc [29] leverages a sparse voxel decoder to reconstruct the sparse geometry of a scene and uses sparse queries to predict masks and labels. OSP [39] treats the scene as a collection of points and uses flexible sampling to allow the model to pay special attention to certain areas. However, all camera-based methods inevitably need to cope with the challenges of adverse lighting and weather conditions, and many works inevitably lack robustness due to the ill-posed monocular depth estimation problem [13, 20].
@@
While camera-based occupancy prediction has demonstrated promising results, multi-modal approaches offer greater reliability and robustness, making them indispensable for practical applications in autonomous driving and robotics. Due to the fact that cameras are susceptible to changing lighting and weather conditions, OccFusion [33] enhances the accuracy and robustness of the occupancy network by integrating features from LiDAR and radar. In concurrent work, OccFusion [58] projects the preprocessed denser and more uniform point cloud onto the image plane to establish a mapping relationship, and performs deformable attention [63] to fuse the corresponding features. Although OccFusion [58] avoids depth estimation, using deformable attention incurs a greater computational burden. Hydra [49] extends FB-BEV [24] to the camera-radar fusion domain and improves the accuracy of depth estimation with the help of radar features. Co-Occ [34] employs K-nearest neighbor search in selecting neighboring camera features to enhance the corresponding LiDAR features and proposes a regularization based on implicit volume rendering. However, this regularization only utilizes the distance ground truth of the point cloud and does not leverage its intrinsic geometry information. EFFOcc [40] proposes an efficient and lightweight multi-modal 3D occupancy network, but it significantly relies on the 3D detection pretraining of the LiDAR branch for optimal results. This dependence, however, limits the flexibility of the network structure design. Specifically, to maximize the benefits of the 3D detection pretraining, EFfOcc has to adjust its network architecture to align more closely with well-established LiDAR detection networks. In addition, the inherently ill-posed nature of the monocular depth estimation problem inevitably renders the image branch of EFFOcc less robust [13]. In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [63, 22] as well as image depth estimation during feature fusion, and does not require 3D object detection pretraining. Furthermore, to fully exploit the inherent geometry information in point cloud features, we add supervision of 3D object detection as an auxiliary branch on the fused features.
@@
Multi-Modal 3D Object Detection: Recent multi-modal 3D object detection methods [30, 18, 19, 5, 43] mainly focus on learning effective BEV feature representations. TransFusion [1] proposes a two-stage transformer-decoder based detection head and applies cross attention to obtain image features for each object query. BEVFusion [30] proposes an efficient and generic multi-task multi-sensor fusion framework which unifies multi-modal features in the shared bird’s-eye view (BEV) representation space, and introduces a specialized kernel to speed up the BEV pooling operation. In concurrent work, to improve the robustness of the LiDAR-camera fusion framework to sensor failures, such as missing LiDAR sensor input, BEVFusion decomposes the LiDAR-camera fusion into two streams that can independently output perception results, and performs feature fusion after the two streams. DAL [13] follows the concept of ‘Detecting As Labeling’ and decouples the fused features during classification and regression. Specifically, it employs the fused features for classification, whereas it utilizes the point cloud features exclusively for regression. In this work, we incorporate the simple yet effective feature fusion approach of BEVFusion and introduce 3D object detection as an auxiliary branch during training.
@@ 3.1 Overall Framework
Our goal is to fully exploit the advantages of point cloud features for multi-modal occupancy prediction. Previous multi-modal studies have not fully exploited this and can only achieve superior performance through more complex image feature extraction networks and larger input image resolutions. The overall framework of our proposed DAOcc is illustrated in Figure 2. DAOcc takes surround images and their corresponding time-synchronized point cloud as input, and obtains the features of the image and point cloud through the Camera Encoder and LiDAR Encoder, respectively. The 2D image features are transformed into the 3D voxel space by projection and sampling. Subsequently, the image and point cloud features in the 3D space are compressed along the height dimension to generate the corresponding BEV features. A simple 2D convolution is then applied for feature fusion, and the fully convolutional BEV Encoder encodes these fused features to obtain the final BEV representations. Finally, the occupancy head uses the Channel-to-Height operation [54] to restore the height of the BEV representations, resulting in final 3D voxel space representations that can be utilized for occupancy prediction. These modules collectively form the basic network architecture of DAOcc, which will be elaborated in Section 3.2.
On the foundation of the basic network, we introduce the BVRE strategy (see Section 3.3) to compensate for the information loss that arises from the reduction in image resolution. This strategy aims to enrich the spatial contextual semantic information by broadening the perceptual range from the Bird’s Eye View (BEV) perspective. Furthermore, to fully leverage the inherent geometric structures within point cloud features, we incorporate 3D object detection as auxiliary supervision (see Section 3.4). This auxiliary supervision not only enhances the discriminability of the fused features but also leads to a very concise overall training loss (see Section 3.5) for our proposed framework.
@@ 3.2 Basic Network
LiDAR Encoder. The method of embedding raw LiDAR points into 3D voxelized features is consistent with BEVFusion [30]. We first voxelize the point cloud, retaining a maximum of 10 points per voxel, resulting in a 3D voxel grid of size $D\times H\times W$.The feature representation for each voxel is obtained by averaging the features of all points within it. Next, we apply 3D sparse convolutions [52] to encode these voxel features, generating spatially compressed LiDAR voxel features $F_{l}\in\mathbb{R}^{C\times\frac{D}{16}\times\frac{H}{8}\times\frac{W}{8}}$, where $C$ represents the feature dimensions.
Camera Encoder. For image feature extraction, taking surround images as input, we first use ResNet50 [11] as the backbone to extract multi-scale features, denoted as $F_{ms}=\{F_{ms\times\frac{1}{8}},F_{\frac{1}{16}},F_{\frac{1}{32}}\}$, where $F_{\frac{1}{x}}$ represent features extracted after $x\times$ downsampling. Then, we employ the Feature Pyramid Network (FPN) [25] as the neck to aggregate these multi-scale features. The output feature map $F_{c_{p}}$ has a shape of $N_{p}\times C_{p}\times\frac{H_{p}}{8}\times\frac{W_{p}}{8}$, where $H_{p}$ and $W_{p}$ represent the input resolution of the image, and $C_{p}$ and $N_{p}$ indicate the number of channels and the number of surround images, respectively.
Projection and Interpolation. For image-related occupancy prediction, a key step is to transform image features from 2D image plane to 3D volume space. Most existing methods use monocular depth estimation [46, 40, 12, 54] or deformable attention [48, 41, 47, 58]. However, monocular depth estimation is inherently an ill-posed problem [13, 20], while deformable attention imposes a significant computational burden [10]. Given these limitations, we use a simple yet effective projection and sample approach similar to Harley et al. [9]. Concretely, we first predefine a 3D voxel grid with a shape of $Z\times\frac{H}{8}\times\frac{W}{8}$, where $Z$ represents the number of voxels along the $z$-axis. The center point of each voxel is then projected onto the image feature plane, and only points that fall within both the image feature plane and the camera’s field of view are retained. Next, the sub-pixel projection positions of the retained points are bilinearly interpolated to generate the image features corresponding to each voxel. For voxels located in the overlapping viewing regions of the surround cameras, we average the image features from the two corresponding cameras to obtain the final feature for each voxel. The output camera voxel features can be denoted as $F_{c}\in\mathbb{R}^{C\times Z\times\frac{H}{8}\times\frac{W}{8}}$.

### v2

@@ Abstract
Multi-sensor fusion significantly enhances the accuracy and robustness of 3D semantic occupancy prediction, which is crucial for autonomous driving and robotics. However, most existing approaches depend on large image resolutions and complex networks to achieve top performance, hindering their application in practical scenarios. Additionally, most multi-sensor fusion approaches focus on improving fusion features while overlooking the exploration of supervision strategies for these features. To this end, we propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image feature extraction network and practical input image resolution. Furthermore, we introduce a BEV View Range Extension strategy to mitigate the adverse effects of reduced image resolution. Experimental results show that DAOcc achieves new state-of-the-art performance on the Occ3D-nuScenes and SurroundOcc benchmarks, and surpasses other methods by a significant margin while using only ResNet50 and 256×704 input image resolution. Code will be made available at https://github.com/AlphaPlusTT/DAOcc.
@@ 1 Introduction
Figure 1: Comparison of loss functions. Unlike existing multi-modal methods [49, 37, 61, 43, 38], which rely on complex combinations of loss functions, including several or all of the focal loss $\mathcal{L}_{focal}$ [41], the scene-class affinity loss $\mathcal{L}_{scal}^{geo}$ and $\mathcal{L}_{scal}^{sem}$ [4], the lovasz-softmax loss $\mathcal{L}_{ls}$ [2], and the depth loss $\mathcal{L}_{d}$ [22], to achieve optimal performance, our approach only requires a single cross-entropy loss $\mathcal{L}_{ce}$ for the occupancy prediction task.
@@
In existing works on multi-modal [60, 37, 61, 52, 43, 49, 38] or image-based [4, 51, 44, 13, 50, 24, 57, 42, 53] occ tasks, achieving superior performance often involves using extremely high image resolutions and complex image feature extraction networks (see Table 1 and Table 3), such as using 900×1600 resolution input image and ResNet101 [12] equipped with DCN [8, 66]. However, this approach significantly limits the deployment of top-performing occ models on edge devices due to the high computational demands. In comparison to images, point cloud is much sparser. For example, in the training set of the nuScenes [3] dataset, the maximum number of points in a single frame point cloud is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image. Therefore, how to effectively leverage point cloud data within a multi-modal occ framework remains to be further explored.
@@
Moreover, we observe that most works [60, 37, 61, 52, 43, 49] on multi-modal occ primarily focuses on obtaining more effective fusion features, with insufficient exploration into the forms of supervision for these fused features. Although CO-Occ [38] introduces regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometry and structure information inherent in point cloud features. In contrast, LiDAR-based 3D detectors [59, 7, 34] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based 3D detectors [28, 17, 29, 30]. This observation suggests a new research direction: how to better exploit the unique strengths of point cloud data in multi-modal occ tasks.
@@
Figure 2: Overview of our proposed DAOcc. We first introduce the BVRE strategy to enrich the spatial contextual information by broadening the perceptual range from the BEV perspective. Then feature extraction is performed through the multi-modal occupancy prediction network illustrated in the above figure. Furthermore, to fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision.
@@
Based on these observations, we propose DAOcc, a novel multi-modal occ framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
@@
In constructing the baseline network of DAOcc, we adopt the simplest and most direct approach. First, we use a 2D image encoder and 3D sparse convolutions [55] to extract features from images and point clouds, respectively. Given that depth estimation from monocular images is an ill-posed problem [14, 21] and that deformable attention modules are overly complex [11], we employ a simple method similar to Harley et al. [10] to transform image features from 2D space to 3D volumetric space. Specifically, we project a set of predefined 3D points onto the 2D image feature plane and use bilinear interpolation to sample the corresponding 2D image features for these points. Next, we adopt a straightforward fusion strategy by concatenating the image and point cloud features and then applying 2D convolution to obtain a unified BEV (Bird’s Eye View) feature. Finally, we apply a fully convolutional BEV encoder with a residual structure to further fuse the unified BEV features, and subsequently employ the Channel-to-Height [57] operation to transform the channel dimension into the height dimension.
@@
To fully exploit the geometry and structure information inherent in point cloud features, we augment the baseline model’s unified BEV features with 3D object detection supervision, thereby enhancing the discriminability of the unified BEV features. This renders the unified BEV features more sensitive to object boundaries and enables our method to avoid the need for complex loss function combinations, as depicted in Figure 1. Additionally, given the sparsity of point clouds, we extend the processing range of the point clouds and employ sparse convolutions [55] to mitigate the computational overhead introduced by this extension. We refer to this approach as BVRE (BEV View Range Extension). BVRE provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution. It is worth noting that the 3D object detection supervision is only used as an auxiliary branch during training and can be removed during inference for the occ task.
@@
Extensive experiments are conducted to validate the efficiency and effectiveness of the proposed DAOcc. The results show that DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes [44] and SurroundOcc [51] benchmarks, while using only ResNet50 [12] and a 256x704 input image resolution. Specifically, on the Occ3D-nuScenes validation set, DAOcc achieves 53.82 mIoU when using the camera mask during training and 48.2 RayIoU without using the camera mask. Additionally, on the SurroundOcc validation set, DAOcc achieves 45.0 IoU and 30.5 mIoU.
@@
  • We design a simple yet efficient multi-modal 3D semantic occupancy prediction baseline, eliminating the need for complex deformable attention [23] modules as well as image depth estimation during feature fusion.
  • We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
@@
  • We establish the new state-of-the-art performance on Occ3D-nuScenes dataset and SurroundOcc dataset, while using only ResNet50 and a 256×704 input image resolution.
@@ 2 Related Work
Camera-based 3D occupancy prediction has gained significant attention due to its cost-effectiveness. Early works [4, 16, 62] lack dense surround occupancy supervision, resulting in sparse occupancy predictions. To obtain dense occupancy prediction, OpenOccupancy [49], SurroundOcc [51] and Occ3D [44] developed methods for generating dense occupancy labels and established benchmarks on their respective proposed datasets. It is worth noting that SurroundOcc does not introduce manual annotation. The emergence of these benchmarks further promoted the development of 3D occupancy prediction. PanoOcc [50] adopts voxel query and a coarse-to-fine approach to learn a unified occupancy representation. FlashOcc [57] uses BEV features with 2D convolution to achieve efficient feature extraction and proposes the Channel-to-Heigh transformation to lift the output of BEV to 3D space. SparseOcc [31] leverages a sparse voxel decoder to reconstruct the sparse geometry of a scene and uses sparse queries to predict masks and labels. However, all camera-based methods inevitably need to cope with the challenges of adverse lighting and weather conditions, and many works inevitably lack robustness due to the ill-posed monocular depth estimation problem [14, 21].
@@
While camera-based occupancy prediction has demonstrated promising results, multi-modal approaches offer greater reliability and robustness, making them indispensable for practical applications in autonomous driving and robotics. Due to the fact that cameras are susceptible to changing lighting and weather conditions, OccFusion [37] enhances the accuracy and robustness of the occupancy network by integrating features from LiDAR and radar. In concurrent work, OccFusion [61] projects the preprocessed denser and more uniform point cloud onto the image plane to establish a mapping relationship, and performs deformable attention [67] to fuse the corresponding features. Although OccFusion [61] avoids depth estimation, using deformable attention incurs a greater computational burden [11]. Hydra [52] extends FB-BEV [25] to the camera-radar fusion domain and improves the accuracy of depth estimation with the help of radar features. Co-Occ [38] employs K-nearest neighbor search in selecting neighboring camera features to enhance the corresponding LiDAR features and proposes a regularization based on implicit volume rendering. However, this regularization only utilizes the distance ground truth of the point cloud and does not leverage its intrinsic geometry information. EFFOcc [43] proposes an efficient and lightweight multi-modal 3D occupancy network, but it significantly relies on the 3D object detection pre-training of the LiDAR branch for optimal results. This dependence, however, limits the flexibility of the network structure design. Specifically, to maximize the benefits of the 3D detection pre-training, EFFOcc has to adjust its network architecture to align more closely with well-established LiDAR detection networks. In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [23] as well as image depth estimation during feature fusion, and does not require 3D object detection pre-training. Furthermore, to fully exploit the inherent geometry information in point cloud features, we add supervision of 3D object detection as an auxiliary supervision on the fused features.
@@
Multi-Modal 3D Object Detection: Recent multi-modal 3D object detection methods [33, 19, 20, 6, 46] mainly focus on learning effective BEV feature representations. TransFusion [1] proposes a two-stage transformer-decoder based detection head and applies cross attention to obtain image features for each object query. BEVFusion [33] proposes an efficient and generic multi-task multi-sensor fusion framework which unifies multi-modal features in the shared BEV representation space, and introduces a specialized kernel to speed up the BEV pooling operation. In concurrent work, BEVFusion [26] decomposes the LiDAR-camera fusion into two streams that can independently output perception results, and performs feature fusion after the two streams. DAL [14] follows the concept of ‘Detecting As Labeling’ and decouples the fused features during classification and regression. Specifically, it employs the fused features for classification, whereas it utilizes the point cloud features exclusively for regression. In this work, we incorporate the simple yet effective feature fusion approach of BEVFusion [33] and introduce 3D object detection as an auxiliary branch during training.
@@ 3.1 Overall Framework
Our purpose is to fully utilize the inherent geometry and structure information in point cloud features within a multi-modal occupancy prediction framework. Previous multi-modal methods have not fully exploited this and can only achieve superior performance through a more complex image encoder and a larger input image resolution. The overall framework of our proposed DAOcc is illustrated in Figure 2. DAOcc takes surround images and their corresponding time-synchronized point cloud as input, and obtains the features of the image and point cloud through the Camera Encoder and LiDAR Encoder, respectively. The 2D image features are transformed into the 3D voxel space by projection and interpolation. Subsequently, the image and point cloud features in the 3D space are compressed along the height dimension to generate the corresponding BEV(Bird’s Eye View) features. A simple 2D convolution is then applied for feature fusion, and the fully convolutional BEV Encoder encodes these fused features to obtain the final BEV representations. Finally, the occupancy head uses the Channel-to-Height operation [57] to restore the height of the BEV representations, resulting in final 3D voxel space representations that can be utilized for occupancy prediction. These modules collectively form the basic network architecture of DAOcc, which will be elaborated in Sec. 3.2.
On the foundation of the basic network, we introduce the BVRE strategy (Sec. 3.3) to compensate for the information loss that arises from the reduction in image resolution. This strategy aims to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective. Furthermore, to fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision (Sec. 3.4). This auxiliary supervision not only enhances the discriminability of the fused features but also leads to a very concise overall training loss (Sec. 3.5) for our proposed framework.
@@ 3.2 Basic Network
LiDAR Encoder. The method of embedding raw LiDAR points into 3D voxelized features is consistent with SECOND [55]. We first voxelize the point cloud, retaining a maximum of 10 points per voxel, resulting in a 3D voxel grid of size $D\times H\times W$.The feature representation for each voxel is obtained by averaging the features of all points within it. Next, we apply 3D sparse convolutions [55] to encode these voxel features, generating spatially compressed LiDAR voxel features $F_{l}\in\mathbb{R}^{C\times\frac{D}{16}\times\frac{H}{8}\times\frac{W}{8}}$, where $C$ represents the feature dimensions.
Camera Encoder. For image feature extraction, taking surround images as input, we first use ResNet50 [12] as the backbone to extract multi-scale features, denoted as $F_{ms}=\{F_{\frac{1}{8}},F_{\frac{1}{16}},F_{\frac{1}{32}}\}$, where $F_{\frac{1}{x}}$ represent features extracted after $x\times$ downsampling. Then, we employ the Feature Pyramid Network (FPN) [27] as the neck to aggregate these multi-scale features. The output feature map $F_{c_{p}}$ has a shape of $N_{p}\times C_{p}\times\frac{H_{p}}{8}\times\frac{W_{p}}{8}$, where $H_{p}$ and $W_{p}$ represent the input resolution of the image, and $C_{p}$ and $N_{p}$ indicate the number of channels and the number of surround images, respectively.
Projection and Interpolation. For image-related occupancy prediction, a key step is to transform image features from 2D image plane to 3D volume space. Most existing methods use monocular depth estimation [49, 43, 13, 57] or deformable attention [51, 44, 50, 61]. However, monocular depth estimation is inherently an ill-posed problem [14, 21], while deformable attention imposes a significant computational burden [11]. Given these limitations, we use a simple yet effective projection and sample approach similar to Harley et al. [10]. Concretely, we first predefine a 3D voxel grid with a shape of $Z\times\frac{H}{8}\times\frac{W}{8}$, where $Z$ represents the number of voxels along the $z$-axis. The center point of each voxel is then projected onto the image feature plane, and only points that fall within both the image feature plane and the camera’s field of view are retained. Next, the sub-pixel projection positions of the retained points are bilinearly interpolated to generate the image features corresponding to each voxel. For voxels located in the overlapping viewing regions of the surround cameras, we average the image features from the two corresponding cameras to obtain the final feature for each voxel. The output camera’s voxel features can be denoted as $F_{c}\in\mathbb{R}^{C\times Z\times\frac{H}{8}\times\frac{W}{8}}$.
@@
Figure 3: Comparison of the impact of resolution increase on computational cost. For LiDAR Encoder, the processing range of the point cloud along the $z$-axis in all experiments is from -5m to 3m. The example in the figure above selects the point cloud with the highest point count in the validation set and aggregates a total of 10 frames.